1. Evaluate each of the following definite integrals. a. 529,-zdy b. Sonata dx2. Integrate. a. sec? tane de b. S tanht dt cosht
Answer #1

a). Given y ye-3 dy subu uz y 23 du= aydy (du/) S. = 1 / Study l yazdy u ln (4) = 1/2 lnlu) u = y²3 S Inly ²-3) dy tc y2-3 9= te tant (4) tanlu) -21- = 1/2 tan (219) Szery da q :to so da tan (2/2) tant (b) 22tu ş Tlu olt FIO TT Streda azq.2a) seco Zano. do Subu uz tano dua secodo Sadu - Intu u= tano - in itano) + secto do tano b) tanht cosht ss secht tankt dt -